OpenStudy (anonymous):

If f(x) = 5x + 4 and g(x) = 6x – 3, what is f(g(x))?

jimthompson5910 (jim_thompson5910):

To find f(g(x)), start with f(x) and replace each x with g(x) like this f(x) = 5x+4 f(g(x)) = 5g(x)+4 Now replace the g(x) on the right side with 6x-3 to get f(g(x))=5(6x-3)+4 and simplify to get f(g(x))=30x-11